OpenStudy (anonymous):

find a unit vector in the direction of w = (-4, 6) I got (-sqrt 13/4 , sqrt 13/ 6) but idk if this is right?

OpenStudy (tkhunny):

\(|w| = \sqrt{(-4)^{2} + (6)^{2}}\) Unit Vector in the same direction? \(\left<\dfrac{-4}{|w|},\dfrac{6}{|w|}\right>\)